# Multi-task learning for molecular electronic structure approaching coupled-cluster accuracy

### v1

@@ IV.3 Evaluating molecular properties
$E^{\rm EGNN}$ $=E_{\rm NN}+2\sum_{n=1}^{n_{e}/2}\epsilon_{n}$ (12)
$\vec{p}^{EGNN}$ $=-2e\sum_{n=1}^{n_{e}/2}\sum_{i\alpha,j\beta}(\tilde{c}_{i,\alpha}^{n})^{*}\tilde{c}_{j,\beta}^{n}\langle\phi_{i,\alpha}|\hat{\vec{r}}|\phi_{j,\beta}\rangle$
$\mathbf{Q}^{\rm EGNN}$ $=2e^{2}\sum_{n=1}^{n_{e}/2}\sum_{i\alpha,j\beta}(\tilde{c}_{i,\alpha}^{n})^{*}\tilde{c}_{j,\beta}^{n}\langle\phi_{i,\alpha}|\hat{\vec{r}}\hat{\vec{r}}|\phi_{j,\beta}\rangle$
$C_{i}^{\rm EGNN}$ $=e\left[Z_{i}-2\sum_{n=1}^{n_{e}/2}\sum_{j\alpha\beta}(\tilde{c}_{i,\alpha}^{n})^{*}\tilde{c}_{j,\beta}^{n}S_{i\alpha,j\beta}\right]$
$B_{ij}^{\rm EGNN}$ $=4\sum_{m,n=1}^{n_{e}/2}\sum_{kl\alpha\beta\gamma\delta}(\tilde{c}_{k,\gamma}^{n})^{*}\tilde{c}_{i,\alpha}^{n}S_{k\gamma,j\beta}(\tilde{c}_{l,\delta}^{m})^{*}\tilde{c}_{j,\beta}^{m}S_{l\delta,i\alpha}$
@@
$\alpha_{0}=2e^{2}\sum_{m=n_{e}/2+1}^{N_{\rm basis}}\sum_{n=1}^{n_{e}/2}\frac{\vec{r}_{mn}\vec{r}_{nm}}{\epsilon_{m}-\epsilon_{n}}$ (14)
where $N_{\rm basis}$ is the number of basis functions of the molecule, and $\vec{r}_{mn}\equiv\sum_{i\alpha,j\beta}(\tilde{c}_{i,\alpha}^{m})^{*}\tilde{c}_{j,\beta}^{n}\langle\phi_{i,\alpha}|\hat{\vec{r}}|\phi_{j,\beta}\rangle$. However, the single-particle approximation used in Eq. (14) does not consider the electric screening effect from electron-electron interaction. We use the EGNN to output a screening matrix $T$ and evaluate the corrected polarizability $\alpha$ as follow:
@@
We evaluate the gas phase standard enthalpy of formation of molecules in Fig. 3 using atomic configurations built by Avogadro [46] and relaxed by the BP86 functional with cc-pVDZ basis set. The total energy at the relaxed atomic configuration is then calculated by our EGNN model. The zero-point energy and thermal vibration, rotation, and translation energy at $T=298.15$ K are also calculated by the BP86 functional with cc-pVDZ basis set implemented in ORCA. Summing all energy terms give the inner energy $U$, and the enthalpy is evaluated as $H\simeq U+Nk_{\rm B}T$ ($N$ is the number of molecules and $k_{\rm B}$ is the Boltzmann constant), where we use the ideal gas law. To obtain the standard enthalpy of formation, we subtract the reference state enthalpy of graphite and hydrogen gas at standard condition. The reference enthalpy for each carbon and hydrogen atom are determined as -38.04672 a.u. and -0.57583 a.u., respectively, using CCSD(T) calculation with cc-pVTZ basis set combined with measured standard enthalpy of formation of atomic carbon, atomic hydrogen, and benzene. Atomic configurations of semiconducting polymers in Fig. 4 are relaxed using the PreFerred Potential v5.0.0 [5, 6].
@@ VII Acknowledgements
$\delta\epsilon_{n}$ $=(\mathbf{c}^{n})^{\dagger}\delta H^{\rm eff}\mathbf{c}^{n}$ (S1)
$\delta\mathbf{c}^{n}$ $=\sum_{m\neq n}\frac{(\mathbf{c}^{m})^{\dagger}\delta H^{\rm eff}\mathbf{c}^{n}}{\epsilon_{n}-\epsilon_{m}}\mathbf{c}_{m}$
@@
$\nabla_{\theta}\epsilon_{n}$ $=(\mathbf{c}^{n})^{\dagger}(\nabla_{\theta}V^{\theta})\mathbf{c}^{n}$ (S2)
$\nabla_{\theta}\mathbf{c}_{n}$ $=\sum_{m\neq n}\frac{(\mathbf{c}^{m})^{\dagger}(\nabla_{\theta}V^{\theta})\mathbf{c}^{n}}{\epsilon_{n}-\epsilon_{m}}\mathbf{c}_{m}$
@@
$\nabla_{\theta}f_{E}$ $=2\sum_{n=1}^{n_{e}/2}\nabla V_{nn}$ (S3)
$\nabla_{\theta}f_{\vec{p}}$ $=-4e\sum_{n=1}^{n_{e}/2}\sum_{m=n_{e}/2}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{mn}}{\epsilon_{n}-\epsilon_{m}}\langle\psi_{n}|\hat{\vec{r}}|\psi_{m}\rangle$
$\nabla_{\theta}f_{\mathbf{Q}}$ $=4e^{2}\sum_{n=1}^{n_{e}/2}\sum_{m=n_{e}/2}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{mn}}{\epsilon_{n}-\epsilon_{m}}\langle\psi_{n}|\hat{\vec{r}}\hat{\vec{r}}|\psi_{m}\rangle$
$\nabla_{\theta}f_{C_{i}}$ $=-4e\sum_{n=1}^{n_{e}/2}\sum_{m=n_{e}/2}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{mn}(\tilde{\mathbf{c}}^{n})^{\dagger}(I_{i}S)\tilde{\mathbf{c}}^{m}}{\epsilon_{n}-\epsilon_{m}}$
$\nabla_{\theta}f_{B_{ij}}$ $=4\sum_{n=1}^{n_{e}/2}\sum_{m=n_{e}/2}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{mn}}{\epsilon_{n}-\epsilon_{m}}$
$\cdot(\tilde{\mathbf{c}}^{n})^{\dagger}(SI_{j}PSI_{i}+SI_{i}PSI_{j})\tilde{\mathbf{c}}^{m}$
where $\nabla V_{mn}\equiv(\tilde{\mathbf{c}}^{m})^{\dagger}\nabla_{\theta}V^{\theta}\tilde{\mathbf{c}}^{n}$, the $N_{\rm basis}\times N_{\rm basis}$ matrix $I_{i}$ is identity in the block diagonal part of atom $i$ and zero elsewhere and $P\equiv 2\sum_{n=1}^{n_{e}/2}\tilde{\mathbf{c}}^{n}(\tilde{\mathbf{c}}^{n})^{\dagger}$. The essential method to avoid numerical instability is to remove terms that can be proved to cancel each other. Taking $\nabla_{\theta}f_{\vec{p}}$ as an example: in Eq. (S2), the summation over $m$ goes through all states except $n$. But as the summed formula is antisymmetric to $m$ and $n$, the terms that $m$ goes from 1 to $n_{e}/2$ cancel each other. Only terms that $m$ goes from $n_{e}/2+1$ to $N_{\rm basis}$ have a non-zero contribution to the final gradient. Therefore, $n$ is always occupied, and $m$ is always unoccupied in the summation. As close-shell molecules always have a finite bandgap, $\epsilon_{n}$ and $\epsilon_{m}$ are not close to each other in any term of the summation, so evaluating Eq. S3 is numerically stable.
Table 1: List of molecule names and number of atomic configurations (superscript) for each molecule in the training and testing dataset.
@@
Table 2: List of the names and serial numbers (superscript, in Ref. [35]) of aromatic molecules in the main text Fig. 3.
@@
Similarly, the gradients of $E_{\rm g}$ and $\alpha$ are as follow:
$\nabla_{\theta}\alpha_{0}$ $=2e^{2}\sum_{m=n_{e}/2+1}^{N_{\rm basis}}\sum_{n=1}^{n_{e}/2}\text{Re}\{\frac{\vec{r}_{mn}\vec{r}_{nm}(\nabla V_{nn}-\nabla V_{mm})}{(\epsilon_{m}-\epsilon_{n})^{2}}$ (S5)
$-2\sum_{k\neq m,n}\frac{\vec{r}_{mn}}{(\epsilon_{m}-\epsilon_{n})}\left[\frac{\vec{r}_{nk}\nabla V_{km}}{(\epsilon_{k}-\epsilon_{m})}+\frac{\vec{r}_{km}\nabla V_{mn}}{(\epsilon_{k}-\epsilon_{n})}\right]\}$
$\nabla_{\theta}f_{\alpha}$ $=(I+\alpha_{0}T)^{-1}(\nabla_{\theta}\alpha_{0})(I-T\alpha)$ (S6)
$-\alpha_{0}(\nabla_{\theta}T)(I+\alpha_{0}T)^{-1}\alpha$
@@
The training domain and out-of-distribution testing testing dataset include 20 and 3 different chemical formula shown as the horizontal axis labels of the first 20 and last 3 columns in Fig. 1c in the main text, respectively. Each chemical formula includes up to 5 different molecules (conformers) taken from the PubChem database. The total number of molecules in the training domain and out-of-distribution testing dataset is 70 and 15, respectively. The full list of molecules and the number of atomic configurations for each molecule are listed in Table. I in this file.
@@
A CCSD(T) calculation with cc-pVTZ basis set is implemented in ORCA [29] for each selected configuration, giving the training labels of total energy, electric dipole and quadrupole moment, Mulliken atomic charge, and Mayer bond order. An EOM-CCSD calculation with cc-pVDZ basis set is then implemented to obtain the first excitation energy (bandgap). Finally, we conduct a polarizability calculation with CCSD and cc-pVDZ basis set. The overlap matrix $S$ and Effective Hamiltonian $F$ is obtained from a BP86 DFT calculations with cc-pVDZ basis set, and a B3LYP calculation is implemented with def2-SVP basis set for comparison.
@@
Aromatic molecules in the main text Fig. 3 include all molecules with serial numbers dividable by 4 and enthalpy of formation provided in Ref. [35]. Details of these aromatic molecules are listed in Table. 2.

### v2

@@ IV.3 Evaluating molecular properties
$E^{\rm EGNN}$ $=E_{\rm NN}+2\sum_{i=1}^{n_{e}/2}\epsilon_{i}$ (13)
$\vec{p}^{EGNN}$ $=-2e\sum_{i=1}^{n_{e}/2}\sum_{I\mu,J\nu}(\tilde{c}_{I,\mu}^{i})^{*}\tilde{c}_{J,\nu}^{i}\langle\phi_{I,\mu}|\hat{\vec{r}}|\phi_{J,\nu}\rangle$
$\mathbf{Q}^{\rm EGNN}$ $=2e^{2}\sum_{i=1}^{n_{e}/2}\sum_{I\mu,J\nu}(\tilde{c}_{I,\mu}^{i})^{*}\tilde{c}_{J,\nu}^{i}\langle\phi_{I,\mu}|\hat{\vec{r}}\hat{\vec{r}}|\phi_{J,\nu}\rangle$
$C_{I}^{\rm EGNN}$ $=e\left[Z_{I}-2\sum_{i=1}^{n_{e}/2}\sum_{J\mu\nu}(\tilde{c}_{I,\mu}^{i})^{*}\tilde{c}_{J,\nu}^{i}S_{I\mu,J\nu}\right]$
$B_{IJ}^{\rm EGNN}$ $=4\sum_{i,j=1}^{n_{e}/2}\sum_{KL\mu\nu\lambda\sigma}(\tilde{c}_{K,\lambda}^{i})^{*}\tilde{c}_{I,\mu}^{i}S_{K\lambda,J\nu}(\tilde{c}_{L,\sigma}^{j})^{*}\tilde{c}_{J,\nu}^{j}S_{L\sigma,I\mu}$
@@
$\alpha_{0}=2e^{2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\sum_{i=1}^{n_{e}/2}\frac{\vec{r}_{ai}\vec{r}_{ia}}{\epsilon_{a}-\epsilon_{i}}$ (15)
where $N_{\rm basis}$ is the number of basis functions of the molecule, and $\vec{r}_{ai}\equiv\sum_{I\mu,J\nu}(\tilde{c}_{I,\mu}^{a})^{*}\tilde{c}_{J,\nu}^{i}\langle\phi_{I,\mu}|\hat{\vec{r}}|\phi_{J,\nu}\rangle$. However, the single-particle approximation used in Eq. (15) does not consider the electric screening effect from electron-electron interaction. We use the EGNN to output a screening matrix $T$ and evaluate the corrected polarizability $\alpha$ as follow:
@@
We evaluate the gas phase standard enthalpy of formation of molecules in Fig. 3 using atomic configurations built by Avogadro [54] and relaxed by the BP86 functional with cc-pVDZ basis set. The total energy at the relaxed atomic configuration is then calculated by our EGNN model. The zero-point energy (ZPE) and thermal vibration, rotation, and translation energy at $T=298.15$ K are also calculated by the BP86 functional with cc-pVDZ basis set implemented in ORCA. The ZPE is corrected by the optimal scaling factor of 1.0393 according to Ref. [55]. Summing all energy terms give the inner energy $U$, and the enthalpy is evaluated as $H\simeq U+Nk_{\rm B}T$ ($N$ is the number of molecules and $k_{\rm B}$ is the Boltzmann constant), where we use the ideal gas law. To obtain the standard enthalpy of formation, we subtract the reference state enthalpy of graphite and hydrogen gas at standard condition. The reference enthalpy for each carbon and hydrogen atom are determined as -38.04639 a.u. and -0.57550 a.u., respectively, using CCSD(T) calculation with cc-pVTZ basis set combined with measured standard enthalpy of formation of atomic carbon, atomic hydrogen, and benzene. Atomic configurations of semiconducting polymers in Fig. 4 are relaxed using the PreFerred Potential v5.0.0 [5, 6].
@@ VII Acknowledgements
$\delta\epsilon_{i}$ $=(\mathbf{c}^{i})^{\dagger}\delta H^{\rm eff}\mathbf{c}^{i}$ (S1)
$\delta\mathbf{c}^{i}$ $=\sum_{p\neq i}\frac{(\mathbf{c}^{p})^{\dagger}\delta H^{\rm eff}\mathbf{c}^{i}}{\epsilon_{i}-\epsilon_{p}}\mathbf{c}^{p}$
@@
$\nabla_{\theta}\epsilon_{i}$ $=(\mathbf{c}^{i})^{\dagger}(\nabla_{\theta}\mathbf{V}^{\theta})\mathbf{c}^{i}$ (S2)
$\nabla_{\theta}\mathbf{c}^{i}$ $=\sum_{p\neq i}\frac{(\mathbf{c}^{p})^{\dagger}(\nabla_{\theta}\mathbf{V}^{\theta})\mathbf{c}^{i}}{\epsilon_{i}-\epsilon_{p}}\mathbf{c}^{p}$
@@
$\nabla_{\theta}f_{E}$ $=2\sum_{i=1}^{n_{e}/2}\nabla V_{ii}$ (S3)
$\nabla_{\theta}f_{\vec{p}}$ $=-4e\sum_{i=1}^{n_{e}/2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{ai}}{\epsilon_{i}-\epsilon_{a}}\langle\psi_{i}|\hat{\vec{r}}|\psi_{a}\rangle$
$\nabla_{\theta}f_{\mathbf{Q}}$ $=4e^{2}\sum_{i=1}^{n_{e}/2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{ai}}{\epsilon_{i}-\epsilon_{a}}\langle\psi_{i}|\hat{\vec{r}}\hat{\vec{r}}|\psi_{a}\rangle$
$\nabla_{\theta}f_{C_{I}}$ $=-4e\sum_{i=1}^{n_{e}/2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{ai}(\tilde{\mathbf{c}}^{i})^{\dagger}(I_{I}S)\tilde{\mathbf{c}}^{a}}{\epsilon_{i}-\epsilon_{a}}$
$\nabla_{\theta}f_{B_{IJ}}$ $=4\sum_{i=1}^{n_{e}/2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\text{Re}\frac{\nabla V_{ai}}{\epsilon_{i}-\epsilon_{a}}$
×$\times(\tilde{\mathbf{c}}^{i})^{\dagger}(SI_{J}PSI_{I}+SI_{I}PSI_{J})\tilde{\mathbf{c}}^{a}$
where $\nabla V_{ai}\equiv(\tilde{\mathbf{c}}^{a})^{\dagger}(\nabla_{\theta}\mathbf{V}^{\theta})\tilde{\mathbf{c}}^{i}$, the $N_{\rm basis}\times N_{\rm basis}$ matrix $I_{J}$ is identity in the block diagonal part of atom $J$ and zero elsewhere. Meanwhile, we define $P\equiv 2\sum_{i=1}^{n_{e}/2}\tilde{\mathbf{c}}^{i}(\tilde{\mathbf{c}}^{i})^{\dagger}$. The essential method to avoid numerical instability is to remove terms that can be proved to cancel each other. Taking $\nabla_{\theta}f_{\vec{p}}$ as an example: in Eq. (S2), the summation over $m$ goes through all states except $n$. But as the summed formula is antisymmetric to $m$ and $n$, the terms that $m$ goes from 1 to $n_{e}/2$ cancel each other. Only terms that $m$ goes from $n_{e}/2+1$ to $N_{\rm basis}$ have a non-zero contribution to the final gradient. Therefore, $n$ is always occupied, and $m$ is always unoccupied in the summation. As close-shell molecules always have a finite bandgap, $\epsilon_{n}$ and $\epsilon_{m}$ are not close to each other in any term of the summation, so evaluating Eq. S3 is numerically stable.
Similarly, the gradients of $E_{\rm g}$ and $\alpha$ are as follow:
$\nabla_{\theta}\alpha_{0}$ $=2e^{2}\sum_{a=n_{e}/2+1}^{N_{\rm basis}}\sum_{i=1}^{n_{e}/2}\text{Re}\left\{\frac{\vec{r}_{ai}\vec{r}_{ia}(\nabla V_{ii}-\nabla V_{aa})}{(\epsilon_{a}-\epsilon_{i})^{2}}\right.$ (S5)
$\left.-2\sum_{p\neq a,i}\frac{\vec{r}_{ai}}{(\epsilon_{a}-\epsilon_{i})}\left[\frac{\vec{r}_{ip}\nabla V_{pa}}{(\epsilon_{p}-\epsilon_{a})}+\frac{\vec{r}_{pa}\nabla V_{ai}}{(\epsilon_{p}-\epsilon_{i})}\right]\right\}$
$\nabla_{\theta}f_{\alpha}$ $=(I+\alpha_{0}T)^{-1}(\nabla_{\theta}\alpha_{0})(I-T\alpha)$
$-\alpha_{0}(\nabla_{\theta}T)(I+\alpha_{0}T)^{-1}\alpha$
Table 1: List of molecule names and number of atomic configurations (labelled in the superscript) for each molecule in the training and testing dataset.
@@
Table 2: List of the names and serial numbers (superscript, defined in Ref. [45]) of aromatic molecules in the main text Fig. 3.
@@
The training domain and out-of-distribution testing testing dataset include 20 and 3 different chemical formula shown as the horizontal axis labels of the first 20 and last 3 columns in Fig. 1c in the main text, respectively. Each chemical formula includes up to 5 different molecules (conformers) taken from the PubChem database. The total number of molecules in the training domain and out-of-distribution testing dataset is 70 and 15, respectively. The full list of molecules and the number of atomic configurations for each molecule are listed in Table 1 in this file.
@@
A CCSD(T) calculation with cc-pVTZ basis set is implemented in ORCA [33] for each selected configuration, giving the training labels of total energy, electric dipole and quadrupole moment, Mulliken atomic charge, and Mayer bond order. An EOM-CCSD calculation with cc-pVDZ basis set is then implemented to obtain the first excitation energy (bandgap). Finally, we conduct a polarizability calculation with CCSD and cc-pVDZ basis set. The overlap matrix $S$ and Effective Hamiltonian $F$ is obtained from a BP86 DFT calculations with cc-pVDZ basis set.
For comparison, B3LYP hybrid DFT calculations are implemented with def2-SVP basis set in ORCA. DSD-PBEP86 double-hybrid DFT calculations are implemented with the def2-TZVP basis set in ORCA. The DM21 double-hybrid DFT calculations are implemented with the def2-TZVP basis set in PySCF program package [56] The AIQM1 calculations are implemented in the MLatom platform [40]. For DM21 and AIQM1, the isolated-atom energies of carbon and hydrogen are re-calibrated according to the least-mean-square criterion to give optimal combination energies in our dataset.
@@
Aromatic molecules in the main text Fig. 3 include all molecules with serial numbers dividable by 4 and enthalpy of formation provided in Ref. [45]. Details of these aromatic molecules are listed in Table 2.
S3 Infrared spectrum
In order to evaluate the infrared spectrum, we first implement a B3LYP hybrid DFT calculation with def2-TZVP basis set to obtain the vibrational modes and frequency of a benzene molecule. Then, we generate atomic configurations displaced from the equilibrium configuration along each vibrational mode by a displacement of -0.1, -0.05, 0.05, and 0.1 Å. Our EGNN model is used to evaluate the electric dipole moment at each configuration, and the dipole-moment derivative with respect to each normal coordinate is evaluated by linear regression. The infrared band intensity of fundamental bands are then evaluated following the method in Ref. [57].
As the two combination bands at 1800 - 2000 cm-1 are mainly contributed by $\nu_{10}+\nu_{17}$ and $\nu_{5}+\nu_{17}$ [58], we generate atomic configurations displaced from the equilibrium configuration by displacement vectors of $0.1(\vec{e}_{i}+\vec{e}_{j}),0.1(\vec{e}_{i}-\vec{e}_{j}),0.1(-\vec{e}_{i}+\vec{e}_{j})$, and $0.1(-\vec{e}_{i}-\vec{e}_{j})$, where $(\vec{e}_{i},\vec{e}_{j})$ are the pair of vibrational modes contribution to each combination band. The second-order dipole-moment derivatives with respect to each pair of normal coordinates $\frac{\partial^{2}\vec{p}}{\partial Q_{i}\partial Q_{j}}$ are then obtained by finite difference method. The leading-order anharmonic constants are also evaluated by finite difference method. These parameters are then used to calculate intensity of the combination bands by Fermi’s golden rule. Using the calculated infrared spectrum peak positions and intensity, we add Gaussian broadening to each peak and fit their bandwidth to the experimental spectrum.